Ida Sidha Karya Company is a family-owned company located in the village of Gianyar on the island of Bali in Indonesia. The company produces a handcrafted Balinese musical instrument called a gamelan that is similar to a xylophone. The gamelans are sold for $870. Selected data for the company’s operations last year follow: Units in beginning inventory 0 Units produced 260 Units sold 230 Units in ending inventory 30 Variable costs per unit: Direct materials $ 105 Direct labor $ 325 Variable manufacturing overhead $ 45 Variable selling and administrative $ 15 Fixed costs: Fixed manufacturing overhead $ 65,000 Fixed selling and administrative $ 21,000 The absorption costing income statement prepared by the company’s accountant for last year appears below: Sales $ 200,100 Cost of goods sold 166,750 Gross margin 33,350 Selling and administrative expense 24,450 Net operating income $ 8,900 Required:

1. Determine how much of the ending inventory consists of fixed manufacturing overhead cost deferred in inventory to the next period.
Answer #1

Fixed Manufacturing Overhead Cost Deferred in Inventory = $7500

Calculated as following

Units produced 260

Units sold = 230

Ending Units = 30

Fixed costs: Fixed manufacturing overhead $ 65,000

Fixed overhead rate for Units Produced = Fixed Manufacturing Overhead/Total units produced = $65000/260 = $250

So $250 per unit is included in valuation a unit produced

so as there are 30 Units unsold

so Fixed Manufacturing Overhead Cost Deferred in Inventory = Ending Units X Fixed overhead in one unit

= 30 x $250

= $7500

Answer 2

Ida Sidha Karya Company
Variable Costing Income Statement
Sales = 230*870 $     200,100
Less. Variable Cost
Direct Material = 230*105 $    24,150
Direct Labor = 230*325 $    74,750
Variable overhead = 230*45 $    10,350
Variable Selling overhead = 230*15 $      3,450
Total Variable Cost $     112,700
Contribution $        87,400
Fixed Manufacturing Overhead $    65,000
Fixed Selling and Administrative overhead $    21,000
Total Fixed Overhead $        86,000
Net Income(Loss) $          1,400

FEEL FREE TO ASK ANY CLARIFICATION IF REQUIRED KINDLY PROVIDE FEED BACK BY THUMBS UP IF SATISFIED IT WILL BE HIGHLY APPRECIATED

Best efforts are made to provide best answer still some error may creep in if something such happens kindly let me know i ll correct the error.

THANK YOU